OpenStudy (zzr0ck3r):

If \(f\) is a nonnegative Lebesgue measurable function and \(\{E_n\}_n^\infty\) is a sequence of Lebesgue measurable sets with \(E_1\subset E_2.....\), then \[\large \int_{\cup_{n=1}^{\infty} E_n} f d\lambda = \lim_{n\rightarrow \infty}\int_{E_n} f d\lambda\]

OpenStudy (zzr0ck3r):

@eliassaab please and thank you.

OpenStudy (anonymous):

Please read this http://en.wikipedia.org/wiki/Monotone_convergence_theorem

OpenStudy (anonymous):

Using the above let \[ E=\cup_{n=1}^\infty E_n \\ f_n = f 1_{E_n}\\ f_E= f 1_E \] then \[ f_n \le f_{n+1} \le \cdots \le f_E\\ \int_{E_n} f d\lambda =\int_Ef_n d\lambda Now apply the above and you are done.

OpenStudy (anonymous):

then \[ f_n \le f_{n+1} \le \cdots \le f_E\\ \int_{E_n} f d\lambda =\int_Ef_n d\lambda \]Now apply the above and you are done.

OpenStudy (zzr0ck3r):

is \(f1_E\) the characteristic rep?

OpenStudy (anonymous):

\( f 1_E\) is f on E and 0 outside E

OpenStudy (zzr0ck3r):

Right

OpenStudy (zzr0ck3r):

Sorry but can you walk.me through how you got from the inequality tho the integral

OpenStudy (zzr0ck3r):

f_n is, covering pointwise?

OpenStudy (zzr0ck3r):

Converging*

OpenStudy (anonymous):

Also it it increasing

OpenStudy (anonymous):

Did you read and understand this http://en.wikipedia.org/wiki/Monotone_convergence_theorem

OpenStudy (zzr0ck3r):

So we are using the pointwise version.of mct?

OpenStudy (anonymous):

The mct is only for pointwise convergence

OpenStudy (zzr0ck3r):

ok I had to get off my tablet, this place is impossible with a tablet... Ok we are using the Lebesgue MCT. We have a \(f_n\) which converges to \(f1_E\) in a increasing manner, and \(f1_E\) is the same thing as \(f_E\), so by MCT we have that \(\lim_{n\rightarrow \infty}\int f_nd\lambda=\int f d\lambda\). Am I right so far?

OpenStudy (anonymous):

Yes

OpenStudy (zzr0ck3r):

so the part im confused about, is when do we drop the limit and where do my bounds come from?

OpenStudy (zzr0ck3r):

I imagine they have something to do with each other.

OpenStudy (anonymous):

What bounds are you referring to?

OpenStudy (anonymous):

Actually \[ \lim_{n\rightarrow \infty}\int f_nd\lambda=\int_E f d\lambda \]

OpenStudy (zzr0ck3r):

how do i get from \(\lim_{n\rightarrow \infty}\int f_nd\lambda=\int f d\lambda\) to \(\int_{E_n} f d\lambda =\int_Ef_n d\lambda\)

OpenStudy (zzr0ck3r):

ok and we get the E_n on the left by the way you constructed f_n?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

ok I get it:) thank you so much, I'm going to go try and explain that to someone else, and that should solidify it:)

OpenStudy (anonymous):

That is a good idea.

OpenStudy (zzr0ck3r):

Is this correct?

OpenStudy (anonymous):

\[ f_n \to f \quad \text {pointwise on E } \]

OpenStudy (anonymous):

The rest is OK